# Human-inspired Perspectives: A Survey on AI Long-term Memory

### v1

@@ Abstract
With the rapid advancement of AI systems, their abilities to store, retrieve, and utilize information over the long term - referred to as long-term memory - have become increasingly significant. These capabilities are crucial for enhancing the performance of AI systems across a wide range of tasks. However, there is currently no comprehensive survey that systematically investigates AI’s long-term memory capabilities, formulates a theoretical framework, and inspires the development of next-generation AI long-term memory systems. This paper begins by systematically introducing the mechanisms of human long-term memory, then explores AI long-term memory mechanisms, establishing a mapping between the two. Based on the mapping relationships identified, we extend the current cognitive architectures and propose the Cognitive Architecture of Self-Adaptive Long-term Memory (SALM). SALM provides a theoretical framework for the practice of AI long-term memory and holds potential for guiding the creation of next-generation long-term memory driven AI systems. Finally, we delve into the future directions and application prospects of AI long-term memory.
@@ 1 Introduction
Over the past few decades, AI has demonstrated significant development potential and astonishing capabilities, thanks to the computational power supported by hardware and advanced machine learning algorithms [hochreiter1997long, nilsson2009quest, vaswani2017attention, xi2023agents, wayne2023llm]. As AI evolves and its applications span various industries, endowing it with long-term memory becomes increasingly important. Long-term memory in AI systems can be understood by analogy with human long-term memory, which stores information for a relatively long period [atkinson1968human]. Long-term memory assists AI systems in several ways. First, it helps acquire general understanding, such as integrating learned knowledge to enhance question-answering systems [qa2020karpukhin, xixin2022gmtkbqa, yang2023care, luo2023chatkbqa]. Additionally, it establishes connections between current and past scenarios by storing key features from frames for timely retrieval, benefiting applications like video understanding [cheng2022xmem, mezghan2022memory, shen2023encode, song2023moviechat, he2024ma]. Furthermore, it aids in mastering procedural abilities, such as developing strategy selection capabilities through production rules or reinforcement learning, thus enhancing agents’ adaptability [davis1977production, quinlan1987generating, ceri1991deriving, kaelbling1996reinforcement, laird2008extending, liu2021automated, liu2022feature].
@@
The human memory system offers valuable insights for designing AI long-term memory, as evidenced by recent AI developments that have, to varying degrees, emulated and incorporated structures akin to those in human long-term memory. For instance, some video understanding systems [cheng2022xmem, song2023moviechat] draw upon the Atkinson-Shiffrin Model [atkinson1968human] (a highly influential human memory model, detailed in Sec. 3) to construct hierarchical memory systems. These systems integrate short-term memory, which temporarily stores visual information, with long-term memory, which recalls historical data to enhance task effectiveness. Moreover, cognitive architectures introduced in works such as  [laird2008extending, sumers2023cognitive] adopted pivotal components of human long-term memory: episodic, semantic, and procedural memory (also discussed in Sec. 3). These cognitive architectures leverage the history event flow to form episodic memory, knowledge sources to form semantic memory, and production rules, code or reinforcement learning to form procedural memory, enabling the creation of agents capable of self-learning and experience accumulation. Moreover, other works, although not explicitly grounded in human long-term memory theories, employ methodologies that closely mirror the processing mechanisms of human long-term memory, as explored later in Sec. 4 and Sec. 5. Therefore, the principles underlying human long-term memory present a promising theoretical foundation for AI researchers to develop more advanced AI systems with robust long-term memory capabilities.
@@
Our research identifies a gap in the existing literature, as there are no comprehensive surveys on AI long-term memory that are grounded in human memory theories (Sec. 2). To bridge this gap, this paper provides a survey of AI long-term memory from the perspective of human long-term memory theory, filling the void in the academic community regarding the prototypes, taxonomies, associated challenges, and system design of AI long-term memory. This survey offers researchers with valuable insights for constructing and optimizing AI systems that utilize long-term memory. The main content of this survey is outlined as follows.
@@
In Sec. 3, we introduce human memory categorized by stages of information processing [atkinson1968human], focusing on episodic memory, semantic memory, and procedural memory, the three key components of human long-term memory [tulving1972episodic, tulving1973encoding, tulving1985memory]. We then classify AI long-term memory into non-parametric and parametric memory in Sec. 4, based on the form of storage [lewis2020retrieval, mallen2023trust]. We also discuss the processing mechanisms and the challenges associated with them. In Sec. 5, we make an attempt to establish the relationships between AI’s and human’s long-term memory based on the review. Following that, in Sec. 6, we proposed the Cognitive Architecture of Self-Adaptive Long-term Memory (SALM) by integrating theories of AI long-term memory. This framework addresses the limitations of long-term memory modules in current cognitive architectures and has the potential to exceed the adaptability of human long-term memory processing mechanisms [shiffrin1969memory, sweller2003evolution, manns2006evolution, allen2013evolution]. It has the promise of serving as a guiding framework for the next generation of AI systems driven by long-term memory. Finally, in Sec. 7, we introduce practical measurement methods and potential applications for AI long-term memory.
@@
    We conduct a narrative review of the relevant work on long-term memory of both human and AI.
@@ 2 Research Background and Methodologies
Research Background. We conduct a comprehensive analysis of articles from 2015 onward, relevant to AI memory surveys. Our search, conducted on October 7th, 2024, utilized the following search terms: (Review(s) or Survey(s) or Taxonomy(ies)) and Memory(ies) and (AI or Artificial Intelligence or Agent(s) or Deep Learning or Machine Learning or Neural Network(s)), and organized the research subjects, related publishers and preprint servers of these articles, as depicted in Fig. 1. The majority of the reviewed articles focus on computer memory and recurrent neural networks (RNNs). The computer memory domain primarily addresses data storage and retrieval, rather than exploring memory taxonomy from an AI perspective [mittal2018survey, umesh2019survey, vurdelja2020survey, radanliev2021review, qian2021graph, asad2022survey, chen2022resistive, haensch2023compute, snasel2023review, ulidowski2023saving, radhika2023review, kaur2024comprehensive, alaparthi2024survey]. In contrast, RNNs, particularly long short-term memory networks (LSTMs), focus on managing hidden states for storing and utilizing sequential historical information, which is not generalizable across all AI domains [bagherzadeh2019review, ma2019taxonomy, li2021quantitative, su2022recurrent, tan2022performance, khan2023short, ghojogh2023recurrent, tyagi2023comparative]. The survey by Savya et al. [savya2023memory] provides a broader analysis of neural network models, including Transformers [vaswani2017attention] and Neural Turing Machines [graves2014neural], and examines the implications of human memory theories, such as the Atkinson-Shiffrin Model [atkinson1968human], for AI memory. However, they focused on the implementation of memory within specific model architectures, and they did not further categorize long-term memory into types such as episodic, semantic, or procedural memory [tulving1985memory] to assess their relevance to AI. On the other hand, the development of Large Language Models (LLMs) [wayne2023llm] has recently accelerated research on memory in intelligent agents. Zhang et al. [zhang2024survey] categorized the memory types in LLM-based agents into textual and parametric forms, which aligned with the long-term memory characteristic of retaining information for an extended period [atkinson1968human]. However, their work does not explicitly incorporate human memory theories to structure the discussion. In summary, current research on memory in AI faces several limitations:
@@
    Current reviews related to memory, particularly long-term memory, are insufficient to encompass the field of artificial intelligence.
@@
    Current reviews on AI-related memory, particularly long-term memory, lack a framework that incorporates theories of human memory.
Research Methodologies. To address these limitations, we adopt human memory theories as the foundation for structuring our review of AI long-term memory. We begin by reviewing literature on human memory theories and their processing stages, with a focus on cognitive science and neuroscience (Sec. 3). We then explore the hierarchy and processing mechanisms of human long-term memory as prototypes to identify AI-related papers exhibiting similar characteristics. Our approach begins with a review of literature on human memory, focusing on its stages of processing as discussed in cognitive science and neuroscience (Sec. 3). We then examine the hierarchy and processing mechanisms related to human long-term memory (Fig. 2), using these as prototypes to identify AI-related studies that exhibit comparable characteristics. Many AI papers related to long-term memory do not explicitly label “long-term memory” as a topic, instead incorporating it implicitly. For instance, numerous studies on long-term memory retrieval, as discussed in Sec. 4.1.2, may not explicitly reference “memory” in their themes; however, they remain intrinsically linked to the retrieval processes associated with human long-term memory. Therefore, in addition to the search for terms like “memory” or “long-term memory”, our review also examines common AI themes such as “neural networks”, “deep learning”, “reinforcement learning”, and more recent trends like “LLMs” and “RAG” (Retrieval-Augmented Generation), as well as interdisciplinary fields like “cognitive architectures”. We classify and synthesize our findings, connecting them to the processing and classification systems of human long-term memory as outlined in Sec. 4 and Sec. 5, and illustrated in Fig. 4. Additionally, we propose an actionable framework based on our human memory-inspired AI long-term memory theory (Sec. 5) and discussed relevant metrics and applications powered by AI long-term memory (Sec. 7). We identify several representative papers to substantiate each point of discussion.
@@ 3 Long-term Memory in Human Brain
Human long-term memory is a subject of extensive research spanning various fields, including cognitive psychology, neuroscience, and computational neuroscience. Insights into human long-term memory enhance our understanding of brain function for researchers in these domains and also offer valuable guidance for AI researchers in constructing effective AI long-term memory. In this section, we provide a comprehensive overview of human memory’s hierarchy (Sec. 3.1) and processing (Sec. 3.2) that are closely related to long-term memory. The overview of the human memory system is illustrated in Fig. 2.
@@ 3.1 Human Memory Hierarchy
Among these models, Atkinson-Shiffrin model, proposed by Atkinson et al., stands out as a significant milestone in memory research. It is recognized for its foundational and comprehensive approach to explain the hierarchical division of memory and its substantial influence on subsequent research in the field. Consequently, we use this model as the theoretical basis for explaining the hierarchy of human memory. Notably, this model and its derivative theories are frequently referenced in AI memory research [lei2023recagent, cheng2022xmem, song2023moviechat], bridging the fields of human and AI memory studies. The Atkinson-Shiffrin model categorizes the human memory system into three levels: Sensory Register (Sec. 3.1.1), Short-term Store (Working Memory)(Sec. 3.1.2), and Long-term Store (Long-term Memory) (Sec. 3.1.3).
@@ 3.1.1 Sensory Register
The sensory register is responsible for receiving and temporarily storing information from sensory systems, allowing visual information to be retained in a highly precise form for a short period. Atkinson et al. [atkinson1968human] used Sperling et al.’s 1960 visual exposure experiment [sperling1960information] to illustrate the existence of the sensory register. In Sperling et al.’s experiment, subjects were briefly exposed to a series of printed letters and then asked to recite them after exposure. They discovered that for exposures lasting between 15 to 500 milliseconds, subjects could correctly report an average of slightly over four letters. When the stimuli contained four or fewer letters, subjects could report them with almost 100% accuracy. This finding demonstrated that individuals can retain brief visual stimuli through a mechanism that reflects the function of the sensory register. Information stored in the sensory register rapidly decays and either vanishes or is transferred to short-term memory within a very short time, providing raw material for subsequent memory processing [atkinson1968human]. This underscores the sensory register’s critical role in the initial stages of information processing and memory formation.
@@ 3.1.2 Working Memory
However, the short-term storage theory proposed by Atkinson et al. has limitations in explaining complex human information processing, particularly regarding how information is processed and stored within the short-term store. They did not systematically introduce the division and details of information processing in the short-term store. Addressing these limitations, Baddeley et al. [baddeley1974working] proposed the working memory theory in 1974, suggesting that working memory is used for the temporary storage and manipulation of information. For example, when performing mental arithmetic, one must not only store numbers temporarily but also apply arithmetic operations to them sequentially. Another example is that recalling numerical information from an oral narrative and completing a reasoning task both require storing and manipulating the information within working memory. Similarly, taking language comprehension to exemplify working memory, individuals must hold words and phrases in memory while simultaneously integrating them into larger syntactic structures and deriving meaning. The manipulations in these examples are essential for tasks like problem-solving, decision-making, and reasoning, which rely on the dynamic handling of information within working memory [baddeley1974working]. From 1974 to 2000, a series of work of Baddeley et al. [baddeley1974working, baddeley1986working, baddeley2000episodic] expanded the working memory theory, dividing it into four components: the Central Executive, Phonological Loop, Visuospatial Sketchpad, and Episodic Buffer. We explain these components briefly while demonstrating them in Fig. 2:
@@
    Central Executive. Baddeley et al. [baddeley1974working, baddeley1986working] highlighted in their work in 1974 and 1986 that the central executive is a critical component of working memory, designed to control and supervise the operation of various modules within working memory, playing a pivotal role in resource allocation similar to that of a CPU (Central Processing Unit) in a computer. When individuals need to process complex transient information, such as temporarily remembering a long sequence of phone numbers, the central executive directs the individual to concentrate more on this task.
@@
    Phonological Loop. In 1986, Baddeley et al. [baddeley1986working] further categorized the phonological loop as a type of working memory, aiming to represent humans’ transient memory for verbal information. The phonemic loop refers to the looping of phonemes in the brain, like an “inner ear” that holds information in a speech-based form for a short time. It is responsible for storing words we hear and silently repeat. The phonemic buffer temporarily stores phonological information, such as phrases, sentences, and numbers. This buffer acts like an “inner voice”, enabling us to repeat information in a loop to maintain and refresh it, preventing it from fading. A demonstration of the phonological loop in action is the process of memorizing a string of phone number by reciting it repeatedly (and silently) until we are able to jot it down in a notebook. In 1998, Baddeley et al. [baddeley1998phonological] reviewed studies on word learning and discovered that typical adults and children use the phonological loop to retain and rehearse the phonological information of words, which aids in the acquisition of new words. However, patients with deficiencies in the phonological loop often have no difficulty in normal language comprehension and spontaneous speech in daily life. This finding suggests that the phonological loop can facilitate the memorization of language information but is not strongly correlated with language proficiency.
@@
    Visuospatial Sketchpad. According to Baddeley et al. [baddeley1986working], the visuospatial sketchpad temporarily stores and processes visual and spatial information. It allows humans to visualize objects and their locations, navigate environments, and plan movements. It integrates visual and spatial information from perception with information stored in long-term memory. This helps in recognizing objects, understanding scenes, and performing tasks that require spatial awareness. This cognitive function is crucial for everyday activities, such as driving, reading maps, engaging in sports, and visualizing and manipulating objects in mind. In 1994, the theory proposed by Kosslyn et al. [kosslyn1994image] provided an enhanced understanding of visuospatial sketchpad. This theory delineates that visuospatial sketchpad can be divided it into three steps: image generation, image maintenance, and image rotation. This process is analogous to solving geometry problems that involve rotating shapes. First, an individual creates an initial mental image of the shape. Next, they maintain the image in their mind. Finally, they imagine how the shape would rotate to match the target image. These researches highlight the critical role of the visuospatial sketchpad, demonstrating its importance in cognitive tasks that require mental manipulation of visual information.
@@
    Episodic Buffer. In 2000, Baddeley et al. [baddeley2000episodic] introduced the episodic buffer as a component of working memory. Episodic buffer facilitates the transfer of information between working memory and long-term memory. Specifically, the episodic buffer enables the movement of significant information from working memory to long-term memory and allows fragments of long-term memory that are relevant to the current context to be retrieved and placed into the episodic buffer. This process supports the formation of new long-term memory and the effective use of long-term memory.

### v2

@@ Abstract
With the rapid advancement of AI systems, their abilities to store, retrieve, and utilize information over the long term - referred to as long-term memory - have become increasingly significant. These capabilities are crucial for enhancing the performance of AI systems across a wide range of tasks. However, there is currently no comprehensive survey that systematically investigates AI’s long-term memory capabilities, formulates a theoretical framework, and inspires the development of next-generation AI long-term memory systems. This paper begins by introducing the mechanisms of human long-term memory, then explores AI long-term memory mechanisms, establishing a mapping between the two. Based on the mapping relationships identified, we extend the current cognitive architectures and propose the Cognitive Architecture of Self-Adaptive Long-term Memory (SALM). SALM provides a theoretical framework for the practice of AI long-term memory and holds potential for guiding the creation of next-generation long-term memory driven AI systems. Finally, we delve into the future directions and application prospects of AI long-term memory.
@@ 1 Introduction
Over the past few decades, AI has demonstrated significant development potential and astonishing capabilities, thanks to the computational power supported by hardware and advanced machine learning algorithms [hochreiter1997long, nilsson2009quest, vaswani2017attention, xi2023agents, wayne2023llm]. As AI evolves and its applications span various industries, endowing it with long-term memory becomes increasingly important. Long-term memory in AI systems can be understood by analogy with human long-term memory, which stores information over periods ranging from days to years [atkinson1968human]. Long-term memory assists AI systems in several ways. First, it helps acquire general understanding, such as integrating learned knowledge to enhance question-answering systems [qa2020karpukhin, xixin2022gmtkbqa, yang2023care, luo2023chatkbqa]. Additionally, it establishes connections between current and past scenarios by storing key features from frames for timely retrieval, benefiting applications like video understanding [cheng2022xmem, mezghan2022memory, shen2023encode, song2023moviechat, he2024ma]. Furthermore, it aids in mastering procedural abilities, such as developing strategy selection capabilities through production rules or reinforcement learning, thus enhancing agents’ adaptability [davis1977production, quinlan1987generating, ceri1991deriving, kaelbling1996reinforcement, laird2008extending, liu2021automated, liu2022feature].
@@
The human memory system offers valuable insights for designing AI long-term memory, as evidenced by recent AI developments that have, to varying degrees, emulated and incorporated structures akin to those in human long-term memory. For instance, some video understanding systems [cheng2022xmem, song2023moviechat] draw upon the Atkinson-Shiffrin Model [atkinson1968human] (a highly influential human memory model, detailed in Sec. 3) to construct hierarchical memory systems. These systems integrate short-term memory, which temporarily stores visual information, with long-term memory, which recalls historical data to enhance task effectiveness. Moreover, cognitive architectures introduced in works such as  [laird2008extending, sumers2023cognitive] adopted pivotal components of human long-term memory: episodic, semantic, and procedural memory (also discussed in Sec. 3). These cognitive architectures leverage the history event flow to form episodic memory, knowledge sources to form semantic memory, and production rules, code or reinforcement learning to form procedural memory, enabling the creation of agents capable of self-learning and experience accumulation. Moreover, other works, although not explicitly grounded in human long-term memory theories, employ methodologies that closely mirror the processing mechanisms of human long-term memory, as explored later in Sec. 4 and Sec. 5. Therefore, the principles underlying human long-term memory present a promising theoretical foundation for AI researchers to develop more advanced AI systems with robust and flexible long-term memory capabilities.
@@
Our research identifies a gap in the existing literature, as there are no comprehensive surveys on AI long-term memory that are grounded in human memory theories (Sec. 2). To bridge this gap, this paper provides a survey of AI long-term memory from the perspective of human memory theory, to motivate and constrain prototypes, taxonomies, associated challenges, and system design of AI long-term memory. This survey offers researchers with valuable insights for constructing and optimizing AI systems that utilize long-term memory. The main content of this survey is outlined as follows.
@@
In Sec. 3, we introduce human memory categorized by stages of information processing [atkinson1968human], focusing on episodic memory, semantic memory, and procedural memory, the three key subsystems of human long-term memory [tulving1972episodic, tulving1973encoding, tulving1985memory]. We then classify AI long-term memory into parametric and non-parametric memory in Sec. 4, based on the form of storage [lewis2020retrieval, mallen2023trust]. We also discuss the processing mechanisms and the challenges associated with these different formats. In Sec. 5, we make an attempt to establish the relationships between AI’s and human’s long-term memory based on the review. Based on these relationships, in Sec. 6, we proposed the Cognitive Architecture of Self-Adaptive Long-term Memory (SALM) by integrating theories of AI long-term memory. This framework addresses the limitations of long-term memory modules in current cognitive architectures and has the potential to exceed the adaptability of human long-term memory processing mechanisms [shiffrin1969memory, sweller2003evolution, manns2006evolution, allen2013evolution]. It has the promise of serving as a guiding framework for the next generation of AI systems driven by long-term memory. Finally, in Sec. 7, we introduce practical measurement methods and potential applications for AI long-term memory.
@@
    We conduct a narrative review of the convergent work on long-term memory of both human and AI.
@@ 2 Research Background and Methodologies
Research Background. We conduct a comprehensive analysis of articles from 2015 onward, relevant to AI memory surveys. Our search, conducted on October 7th, 2024, utilized the following search terms: (Review(s) or Survey(s) or Taxonomy(ies)) and Memory(ies) and (AI or Artificial Intelligence or Agent(s) or Deep Learning or Machine Learning or Neural Network(s)), and organized the research subjects, related publishers and preprint servers of these articles, as depicted in Fig. 1. The majority of the reviewed articles focus on computer memory and recurrent neural networks (RNNs). The computer memory domain primarily addresses data storage and retrieval, rather than exploring memory taxonomy from an AI perspective [mittal2018survey, umesh2019survey, vurdelja2020survey, radanliev2021review, qian2021graph, asad2022survey, chen2022resistive, haensch2023compute, snasel2023review, ulidowski2023saving, radhika2023review, kaur2024comprehensive, alaparthi2024survey]. In contrast, RNNs, particularly long short-term memory networks (LSTMs), focus on managing hidden states for storing and utilizing sequential historical information, which is not generalizable across all AI domains [bagherzadeh2019review, ma2019taxonomy, li2021quantitative, su2022recurrent, tan2022performance, khan2023short, ghojogh2023recurrent, tyagi2023comparative]. The survey by Savya et al. [savya2023memory] provides a broader analysis of neural network models, including Transformers [vaswani2017attention] and Neural Turing Machines [graves2014neural], and examines the implications of human memory theories, such as the Atkinson-Shiffrin Model [atkinson1968human], for AI memory. However, they focused on the implementation of memory within specific model architectures, and they did not further categorize long-term memory into types such as episodic, semantic, or procedural memory [tulving1985memory] to assess their relevance to AI. On the other hand, the development of Large Language Models (LLMs) [wayne2023llm] has recently accelerated research on memory in intelligent agents. Zhang et al. [zhang2024survey] categorized the memory types in LLM-based agents into textual and parametric forms, which aligned with the long-term memory characteristic of retaining information for an extended period [atkinson1968human]. However, their work does not explicitly incorporate human memory theories to structure the discussion. In summary, current surveys on memory in the field of AI, particularly long-term memory, face several limitations:
@@
    Lack a focus on AI systems as a whole.
@@
    Lack a framework that incorporates theories of human memory.
Research Methodologies. To address these limitations, we adopt human memory theories as the foundation for structuring our review of AI long-term memory. Our approach begins with a review of literature on human memory, focusing on its stages of processing as discussed in cognitive science and neuroscience (Sec. 3). We then examine the hierarchy and processing mechanisms related to human long-term memory (Fig. 2), using these as prototypes to identify AI-related studies that exhibit comparable characteristics (Sec. 4 and Sec. 5). Many AI papers related to long-term memory do not explicitly label “long-term memory” as a topic, instead incorporating it implicitly. For instance, numerous studies on long-term memory retrieval, as discussed in Sec. 4.1.2, may not explicitly reference “memory” in their themes; however, they remain intrinsically linked to the retrieval processes associated with human long-term memory. Therefore, in addition to the search for terms like “memory” or “long-term memory”, our review also examines common AI themes such as “neural networks”, “deep learning”, “reinforcement learning”, and more recent trends like “LLMs” and “RAG” (Retrieval-Augmented Generation), as well as interdisciplinary fields like “cognitive architectures”. We classify and synthesize our findings, connecting them to the processing and classification systems of human long-term memory as outlined in Sec. 4 and Sec. 5, and illustrated in Fig. 4. Additionally, we propose an actionable framework based on our human memory-inspired AI long-term memory theory (Sec. 6) and discussed relevant metrics and applications powered by AI long-term memory (Sec. 7). We identify several representative papers to substantiate each point of discussion.
@@ 3 Long-term Memory in Human Brain
Human long-term memory is a subject of extensive research spanning various fields, including cognitive psychology, neuroscience, and computational neuroscience. Insights into human long-term memory enhance our understanding of brain function for researchers in these domains and also offer valuable guidance for AI researchers in constructing effective AI long-term memory. In this section, we provide an overview of human memory’s hierarchy (Sec. 3.1) and processing (Sec. 3.2) that are closely related to long-term memory. The overview of the human memory system is illustrated in Fig. 2.
@@ 3.1 Human Memory Hierarchy
Among these models, Atkinson-Shiffrin model, proposed by Atkinson et al., represents a significant milestone in memory research. It is recognized for its foundational and comprehensive approach to explain the hierarchical division of memory and its substantial influence on subsequent research in the field. Consequently, we use this model to exemplify the theoretical basis for explaining the hierarchy of human memory. Notably, this model and its derivative theories are frequently referenced in AI memory research [lei2023recagent, cheng2022xmem, song2023moviechat], bridging the fields of human and AI memory studies. The Atkinson-Shiffrin model categorizes the human memory system into three levels: Sensory Register (Sec. 3.1.1), Short-term Store (Working Memory)(Sec. 3.1.2), and Long-term Store (Long-term Memory) (Sec. 3.1.3).
@@ 3.1.1 Sensory Register
According to Atkinson et al. [atkinson1968human], the sensory register is responsible for receiving and temporarily storing information from sensory systems, allowing information like visual stimuli to be retained in a highly precise form for a short period. They used Sperling et al.’s 1960 visual exposure experiment [sperling1960information] to illustrate the existence of the sensory register. In Sperling et al.’s experiment, subjects were briefly exposed to a series of printed letters and then asked to recite them after exposure. They discovered that for exposures lasting between 15 to 500 milliseconds, subjects could correctly report an average of slightly over four letters. When the stimuli contained four or fewer letters, subjects could report them with almost 100% accuracy. This finding demonstrated that individuals can retain brief visual stimuli through a mechanism that reflects the function of the sensory register. Information stored in the sensory register rapidly decays and either vanishes or is transferred to short-term memory within a very short time, providing raw material for subsequent memory processing [atkinson1968human]. This underscores the sensory register’s critical role in the initial stages of information processing and memory formation.
@@ 3.1.2 Working Memory
However, the short-term storage theory proposed by Atkinson et al. has limitations in explaining complex human information processing, particularly regarding how information is processed and stored within the short-term store. They did not systematically introduce the division and details of information processing in the short-term store. Addressing these limitations, Baddeley et al. [baddeley1974working] proposed the working memory theory in 1974, suggesting that working memory is used for the temporary storage and manipulation of information. For example, when performing mental arithmetic, one must not only store numbers temporarily but also apply arithmetic operations to them sequentially. Another example is that recalling numerical information from an oral narrative and completing a reasoning task both require storing and manipulating the information within working memory. Similarly, taking language comprehension to exemplify working memory, individuals must hold words and phrases in memory while simultaneously integrating them into larger syntactic structures and deriving meaning. The manipulations in these examples are essential for tasks like problem-solving, decision-making, and reasoning, which rely on the dynamic handling of information within working memory [baddeley1974working]. From 1974 to 2000, a series of works by Baddeley et al. [baddeley1974working, baddeley1986working, baddeley2000episodic] expanded the working memory theory, dividing it into four components: the Central Executive, Phonological Loop, Visuospatial Sketchpad, and Episodic Buffer. We explain these components briefly while summarizing them in Fig. 2:
@@
    Central Executive. Baddeley et al. [baddeley1974working, baddeley1986working] highlighted that the central executive is a critical component of working memory, designed to control and supervise the operation of various modules within working memory, playing a pivotal role in resource allocation similar to that of a CPU (Central Processing Unit) in a computer. When individuals need to process complex transient information, such as temporarily remembering a long sequence of phone numbers, the central executive directs the individual to concentrate more on this task.
@@
    Phonological Loop. Baddeley et al. [baddeley1986working] further categorized the phonological loop as a type of working memory, aiming to represent humans’ transient memory for verbal information. The phonemic loop refers to the looping of phonemes in the brain, like an “inner ear” that holds information in a speech-based form for a short time. It is responsible for storing words we hear and silently repeat. The phonemic buffer temporarily stores phonological information, such as phrases, sentences, and numbers. This buffer acts like an “inner voice”, enabling us to repeat information in a loop to maintain and refresh it, preventing it from fading. A demonstration of the phonological loop in action is the process of memorizing a string of phone number by reciting it repeatedly (and silently) until we are able to jot it down in a notebook. The phonological loop plays a crucial role in memorizing linguistic information, such as new words [baddeley1998phonological].
@@
    Visuospatial Sketchpad. According to Baddeley et al. [baddeley1986working], the visuospatial sketchpad temporarily stores and processes visual and spatial information. It integrates visual and spatial information from perception with information stored in long-term memory. This helps in recognizing objects, understanding scenes, and performing tasks that require spatial awareness, such as driving, reading maps, engaging in sports, and visualizing and manipulating objects in mind. In 1994, the theory proposed by Kosslyn et al. [kosslyn1994image] provided an enhanced understanding of visuospatial sketchpad. This theory delineates that visuospatial sketchpad can be divided it into three steps: image generation, image maintenance, and image rotation. This process is analogous to solving geometry problems that involve rotating shapes. First, an individual creates an initial mental image of the shape. Next, they maintain the image in their mind. Finally, they imagine how the shape would rotate to match the target image. These researches highlight the critical role of the visuospatial sketchpad, demonstrating its importance in cognitive tasks that require mental manipulation of visual information.
@@
    Episodic Buffer. Baddeley et al. [baddeley2000episodic] introduced the episodic buffer as a component of working memory. Episodic buffer facilitates the transfer of information between working memory and long-term memory. Specifically, the episodic buffer enables the movement of significant information from working memory to long-term memory and allows fragments of long-term memory that are relevant to the current context to be retrieved and placed into the episodic buffer. This process supports the formation of new long-term memory and the effective use of long-term memory.